OpenStudy (anonymous):

Elimination. 15x - 10y = 200 -15x - 15y = 1350/17 -25y = 4750/17 how are they getting 4750/17 ???

OpenStudy (anonymous):

Add both 15x-10y-15x-15y=200+(1350/17) -25y=4750/17

OpenStudy (anonymous):

200 + 1350/17 = 200/1 + 1350/17 = 3400/17 + 1350/17 common denominator = 4750/17 add numerators

OpenStudy (anonymous):

Thank you all